It takes the pendulum 3 seconds to swing from A through B to C and back to B. Calculate the periodic time of the pendulum.


1
Expert's answer
2021-05-27T08:30:59-0400

Let time taken to reach A to B = x

Time taken to reach B to C = x

Time taken to reach C to B again = x

Total time = 3 s

"x+x+x=3"

"x=1\space s"

Total time taken to reach point A again "=3+1=4\space s"

therefore time period = 4 s